# Attention as an Adaptive Filter

### v2

@@ 3.1 Generalizing the Adaptive Filter to a Tensor Form of Attention
We now generalize the batch MLE form of the above adaptive filter using learned, complex-valued query, key, value, and output matrices: $\mathbf{W_{q}}\in\mathbb{C}^{d_{qk}\times d_{e}}$, $\mathbf{W_{k}}\in\mathbb{C}^{d_{qk}\times d_{e}}$, $\mathbf{W_{v}}\in\mathbb{C}^{d_{v}\times d_{e}}$, $\mathbf{W_{p}}\in\mathbb{C}^{d_{e}\times d_{v}}$, where $d_{e}$ is the input embedding dimension $d_{qk}$ is the query/key embedding dimension, and $d_{v}$ is the value embedding dimension.
In this formulation, $\mathbf{W_{q}}$, $\mathbf{W_{k}}$, $\mathbf{W_{v}}$ replace $\mathbf{S}^{-1}$ and $\mathbf{W_{p}}$ replaces $\mathbf{S}$. When $\mathbf{W_{q}}=\mathbf{W_{k}}=\mathbf{W_{v}}=\mathbf{W_{p}}^{-1}$, the attention layer corresponds to an adaptive MLE. By allowing these parameter matrices to vary freely, the network can learn approximate inverses as needed, enabling attention to implement an MLE while retaining the flexibility to represent richer transformations.
We can define separate key/query and value dynamics parameters: $\bm{\lambda}_{\mathbf{qk}},\bm{\lambda}_{Q\mathbf{qk}},\bm{\lambda}_{R\mathbf{qk}},\bm{\lambda}_{C\mathbf{qk}}\in\mathbb{R}^{d_{qk}}$; $\bm{\lambda}_{\mathbf{v}},\bm{\lambda}_{Q\mathbf{v}},\bm{\lambda}_{R\mathbf{v}},\bm{\lambda}_{C\mathbf{v}}\in\mathbb{R}^{d_{v}}$. If we set $d_{qk}=d_{v}$, then we can use a shared set of parameters for the key and value dynamics: $\bm{\lambda},\bm{\lambda}_{Q},\bm{\lambda}_{R},\bm{\lambda}_{C}\in\mathbb{R}^{d_{v}}$. Note that we can absorb $\big|\bm{\lambda}_{C}\big|^{2}$ into $\bm{\lambda}_{Q}$, but we keep it here for clarity.
@@
The key and value precision tensors: $\mathcal{P}^{C}_{\mathbf{qk}}\in\mathbb{R}^{d_{qk}\times m\times m},\quad\mathcal{P}^{C}_{\mathbf{v}}\in\mathbb{R}^{d_{v}\times m\times m}$:
$\mathcal{P}^{C}_{\mathbf{qk}}[:,i,j]=\left(\bm{\lambda}_{C\mathbf{qk}}^{2}\odot\bm{\lambda}_{Q\mathbf{qk}}\odot\frac{\mathbf{1}-e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})|t_{j}-t_{i}|}}{-2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})}+\bm{\lambda}_{R\mathbf{qk}}\odot e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})|t_{j}-t_{i}|}\right)^{-1}$
$\mathcal{P}^{C}_{\mathbf{v}}[:,i,j]=\left(\bm{\lambda}_{C\mathbf{v}}^{2}\odot\bm{\lambda}_{Q\mathbf{v}}\odot\frac{\mathbf{1}-e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{v}})|t_{j}-t_{i}|}}{-2\mathrm{Re}(\bm{\lambda}_{\mathbf{v}})}+\bm{\lambda}_{R\mathbf{v}}\odot e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{v}})|t_{j}-t_{i}|}\right)^{-1}$
We apply a causal attention mask to $\mathcal{P}^{C}_{\mathbf{qk}},\mathcal{P}^{C}_{\mathbf{v}}$, masking out all $j>i$.
@@
$\mathbf{W}[i,j]=\left(1+\sum_{k}\mathcal{P}^{C}_{\mathbf{qk}}[k,i,j]\cdot\big|\mathcal{R}_{\mathbf{qk}}[k,i,j]\big|^{2}/\nu\right)^{-1}$
@@
$\tilde{\mathcal{A}}[k,i,j]=\mathbf{W}[i,j]\cdot\mathcal{P}_{\mathbf{v}}^{C}[k,i,j]$
@@ 3.2 Complexity
Vanilla attention with input length $m$ and embedding dimension $d$ has time complexity $\mathcal{O}(m^{2}d)$ to compute attention scores and apply them across the sequence, plus $\mathcal{O}(md^{2})$ to transform the inputs into query, key, and value vectors and to transform the outputs through a linear layer. The storage of the query, key, and value matrices requires $\mathcal{O}(md)$ memory, and the attention matrix requires $\mathcal{O}(m^{2})$ memory.
The naive tensor implementation of attention above involves element-wise multiplication of $d\times m\times m$ tensors, resulting in time complexity $\mathcal{O}(m^{2}d+md^{2})$, which is the same as vanilla attention, but slower in practice due to less-optimized element-wise operations. The time complexity to propagate the precision matrices $\mathbf{P}_{ij}^{C}$ through the differential Lyapunov equation in the full (non-diagonalizable case) is $\mathcal{O}(d^{3})$, plus another $\mathcal{O}(d^{3})$ to invert each of these matrices, for each of the $\mathcal{O}(m^{2})$ time differences. Hence, the total time complexity is $\mathcal{O}(m^{2}d^{3})$. In contrast, for systems that admit a closed-form solution, the propagation and matrix inversions become $\mathcal{O}(d)$, so the total time complexity is $\mathcal{O}(m^{2}d)$. If the measurements are sampled at equal time intervals, the time complexity reduces to $\mathcal{O}(md)$.
@@ 3.3 Convolutional Representation
$\mathcal{K}_{\mathbf{qk}}^{V}[\tau]=\bm{\lambda}_{C\mathbf{qk}}^{2}\odot\bm{\lambda}_{Q\mathbf{qk}}\odot\frac{\mathbf{1}-e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})\Delta t\tau}}{-2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})}+\bm{\lambda}_{R\mathbf{qk}}\odot e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})\Delta t\tau}$
@@
$\mathcal{K}_{\mathbf{qk}}^{V}[\tau]=\bm{\lambda}_{C\mathbf{qk}}^{2}\odot\bm{\lambda}_{Q\mathbf{qk}}\odot\frac{\mathbf{1}-\mathcal{K}^{\dagger}_{\mathbf{k}}[\tau]\mathcal{K}_{\mathbf{qk}}[\tau]}{-2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})}+\bm{\lambda}_{R\mathbf{qk}}\odot\big(\mathcal{K}^{\dagger}_{\mathbf{k}}[\tau]\mathcal{K}_{\mathbf{qk}}[\tau]\big)$
@@
$\mathcal{P}_{\mathbf{k}}^{C}[k,i,j]=\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|],$
where $\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|]=\mathcal{K}_{\mathbf{qk}}^{V}[k,|i-j|]^{-1}$, and similarly for $\mathcal{K}_{\mathbf{v}}^{P}[\tau]$ and $\mathcal{P}_{\mathbf{v}}^{C}$.
@@
3.4 Broadcasting the Estimates
To avoid explicitly constructing the pulled-forward estimates $\mathcal{\hat{Z}}_{\mathbf{k}}[k,i,j]$ and $\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]$, we can define separate arrays for the $i$ and $j$ components:
@@ 3.5 Factorizing the Residuals
$\mathbf{D}[i,j]:=\sum_{k}\mathcal{R}^{*}_{\mathbf{qk}}[k,i,j]\cdot\mathcal{P}^{C}_{\mathbf{qk}}[k,i,j]\cdot\mathcal{R}_{\mathbf{qk}}[k,i,j]$
@@
$\sum_{k}\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\left(\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]-\mathbf{\hat{Z}}^{*}_{\mathbf{k}}[k,j]\right)\cdot\mathcal{P}^{C}_{\mathbf{qk}}[k,i,j]\cdot\left(\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]-\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)$
@@
$\mathbf{D}[i,j]=\sum_{k}\mathcal{K}^{C}_{\mathbf{qk}}[k,|i-j|]\cdot\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\cdot\Big[\big|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\big|^{2}+\big|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big|^{2}-2\mathrm{Re}\big(\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big)\Big]$
@@
3.6 Isotropic Decay and Noise
@@
This corresponds to a system with an isotropic plus skew-symmetric state matrix, $\mathbf{A}=\alpha\mathbf{I}+\bm{\Omega}$, where $\bm{\Omega}\in\mathbb{R}^{d\times d}$ is skew-symmetric ($\bm{\Omega}^{\top}=-\bm{\Omega}$) and diagonalizable over $\mathbb{C}$, that is $\bm{\Omega}=\mathbf{S}\bm{\Lambda}_{\Omega}\mathbf{S}^{-1}$, where $\mathbf{S}\in\mathbb{C}^{d\times d}$, $\bm{\Lambda}_{\Omega}=\text{diag}(\omega_{1},\dots\omega_{d})$, with $\omega_{k}\in i\mathbb{R}$. We also assume that the noise is isotropic, i.e. that the noise parameters are scalar multiples of identity: $\bm{\Lambda}_{C}=\lambda_{C}\mathbf{I},\bm{\Lambda}_{Q}=\sigma^{2}\mathbf{I},\bm{\Lambda}_{R}=\eta^{2}\mathbf{I}$.
Note that we could define separate query/key and value dynamic parameters $\alpha$ and $\bm{\omega}$, but here we use shared parameters for simplicity.
Our covariance kernel becomes:
$\mathcal{K}^{V}(\tau)=\sigma_{V}^{2}(\tau)\mathbf{I}_{d},\quad\text{where:}\quad\sigma_{V}^{2}(\tau)=\big|\lambda_{C}\big|^{2}\sigma^{2}\frac{1-e^{-2\alpha\Delta t\tau}}{-2\alpha}+\eta^{2}e^{-2\alpha\Delta t\tau}$
@@
$\mathbf{P}_{ij}^{C}=\frac{1}{\sigma_{V}^{2}(|i-j|)}.$
We can then pull $\mathbf{P}^{C}_{\mathbf{qk}}\in\mathbb{R}^{m\times m}$ outside the sum:
$\mathbf{D}[i,j]=\mathbf{P}^{C}_{\mathbf{qk}}[i,j]\sum_{k}\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\Big[\big|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\big|^{2}+\big|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big|^{2}-2\mathrm{Re}\left(\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)\Big]$
@@
Defining a score matrix $\mathbf{L}=\log(\mathbf{P}_{\mathbf{v}}^{C})-\log\left(\mathbf{I}+\mathbf{P}^{C}_{\mathbf{qk}}\odot\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}/\nu\right)$, and a causal mask $\mathbf{M}_{\text{causal}}$, we can then express the row-normalization using softmax:
@@
Whereas for the query-key dynamics, the isotropic decay condition $\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})=\alpha\mathbf{1}$ is necessary to reduce the memory requirement, there is no such constraint on $\bm{\lambda}_{\mathbf{v}}$. However, we face another problem: for $\mathrm{Re(}\bm{\lambda}_{\mathbf{v}})\leq\mathbf{0}$, $\mathbf{\hat{Z}_{v}}$ grows exponentially and $\bm{\Phi}_{\mathbf{v}}$ decays exponentially. However, if we again assume isotropic decay $\mathrm{Re}(\bm{\lambda}_{\mathbf{v}})=\alpha\mathbf{1}$, we can re-factorize $\mathbf{\bar{Z}_{v}}[k,i]$ such that this problem is avoided.
@@
$\mathbf{Z}_{\mathbf{v}}^{+}=e^{\bm{\lambda}\Delta t}\odot\mathbf{\bar{Z}_{v}}$
This factorization, which separates scalar decay/growth (magnitude) from oscillatory components (direction), reduces the memory requirement of the algorithm from $\mathcal{O}(m^{2}d)$ to $\mathcal{O}(m^{2}+md)$, as in vanilla attention, while maintaining the $\mathcal{O}(m^{2}d)$ compute time. This comes at the price of losing the second-order information provided by the full tensors $\mathcal{P}^{C}_{\mathbf{qk}}$ and $\mathcal{P}^{C}_{\mathbf{v}}$. Whereas $\mathcal{P}^{C}_{\mathbf{qk}},\mathcal{P}^{C}_{\mathbf{v}}$ represent the confidence of the $k$-th feature’s projection from time $t_{j}$ to time $t_{i}$, $\mathbf{P}^{C}_{\mathbf{qk}},\mathbf{P}^{C}_{\mathbf{v}}$ represent measurements of overall confidence between query and key.
Computing the residual matrix $\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}$ requires additional vector summations, broadcasting, and element-wise multiplications, as opposed to dot-product attention’s single matrix product. However, computing $|\mathbf{Z_{q}|}$ and $|\mathbf{Z_{k}|}$ is only $\mathcal{O}(d_{qk}m)$, so it is negligible overhead compared to the $\mathcal{O}(d_{qk}m^{2})$ dot product computation.
The assumption that the real part of the eigenvalues is a shared scalar serves two purposes: (a) it allows the cross term to be represented as a matrix multiplication, which reduces the memory requirement, and (b) it improves the exploding/vanishing gradient problem for $\alpha<0$, because rather than computing $e^{\alpha t_{i}}$ and $e^{-\alpha t_{j}}$ separately, we compute $e^{\alpha(t_{i}-t_{j})}$, which is stable, since $t_{i}>t_{j}$ for causal systems.
[…]
Practically, the differences are:
  • AFA uses complex-valued $\mathbf{W_{q}}$, $\mathbf{W_{k}}$, $\mathbf{W_{v}}$, $\mathbf{W_{p}}$, whereas RetNet uses real-valued projections.
  • AFA also applies the matrix exponential to the value projections, whereas RetNet does not, keeping its values static.
  • AFA incorporates uncertainty priors into attention using propagated covariance kernels.
3.7 Normalized Queries & Keys

### v3

@@ 3.1 Generalizing the Adaptive Filter to a Tensor Form of Attention
We now generalize the above IRLS step using learned, complex-valued query, key, value, and output matrices: $\mathbf{W_{q}}\in\mathbb{C}^{d_{qk}\times d_{e}}$, $\mathbf{W_{k}}\in\mathbb{C}^{d_{qk}\times d_{e}}$, $\mathbf{W_{v}}\in\mathbb{C}^{d_{v}\times d_{e}}$, $\mathbf{W_{p}}\in\mathbb{C}^{d_{e}\times d_{v}}$, where $d_{e}$ is the input embedding dimension $d_{qk}$ is the query/key embedding dimension, and $d_{v}$ is the value embedding dimension.
In this formulation, $\mathbf{W_{q}}$, $\mathbf{W_{k}}$, $\mathbf{W_{v}}$ replace $\mathbf{S}^{-1}$ and $\mathbf{W_{p}}$ replaces $\mathbf{S}$. When $\mathbf{W_{q}}=\mathbf{W_{k}}=\mathbf{W_{v}}=\mathbf{W_{p}}^{-1}$, the attention layer corresponds to an adaptive filter. By allowing these parameter matrices to vary freely, the network can learn approximate inverses as needed, enabling attention to implement an optimal filter while retaining the flexibility to represent richer transformations.
We can define separate dynamics and noise parameters for the queries/keys and the values: $\bm{\lambda}_{\mathbf{qk}},\bm{\lambda}_{Q\mathbf{qk}},\bm{\lambda}_{R\mathbf{q}},\bm{\lambda}_{R\mathbf{k}},\bm{\lambda}_{C\mathbf{qk}}\in\mathbb{R}^{d_{qk}}$; $\bm{\lambda}_{\mathbf{v}},\bm{\lambda}_{Q\mathbf{v}},\bm{\lambda}_{R\mathbf{v}},\bm{\lambda}_{\Gamma\mathbf{v}},\bm{\lambda}_{C\mathbf{v}}\in\mathbb{R}^{d_{v}}$. Alternatively, if we set the embedding dimension of the queries/keys equal to that of the values, $d_{qk}=d_{v}$, then we can use a shared set of parameters: $\bm{\lambda},\bm{\lambda}_{Q},\bm{\lambda}_{R},\bm{\lambda}_{\Gamma},\bm{\lambda}_{C}\in\mathbb{R}^{d_{v}}$. Note that we can absorb $\big|\bm{\lambda}_{C}\big|^{2}$ into $\bm{\lambda}_{Q}$, but we keep it here for clarity.
@@
The key and value precision tensors: $\mathcal{P}^{Z}_{\mathbf{qk}}\in\mathbb{R}^{d_{qk}\times m\times m},\quad\mathcal{P}^{Z}_{\mathbf{v}}\in\mathbb{R}^{d_{v}\times m\times m}$:
$\mathcal{P}^{Z}_{\mathbf{qk}}[:,i,j]=\left(\bm{\lambda}_{C\mathbf{qk}}^{2}\odot\bm{\lambda}_{Q\mathbf{qk}}\odot\frac{\mathbf{1}-e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})|t_{j}-t_{i}|}}{-2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})}+\bm{\lambda}_{R\mathbf{k}}\odot e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})|t_{j}-t_{i}|}+\bm{\lambda}_{R\mathbf{q}}\right)^{-1}$
$\mathcal{P}^{Z}_{\mathbf{v}}[:,i,j]=\left(\bm{\lambda}_{C\mathbf{v}}^{2}\odot\bm{\lambda}_{Q\mathbf{v}}\odot\frac{\mathbf{1}-e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{v}})|t_{j}-t_{i}|}}{-2\mathrm{Re}(\bm{\lambda}_{\mathbf{v}})}+\bm{\lambda}_{R\mathbf{v}}\odot e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{v}})|t_{j}-t_{i}|}+\bm{\lambda}_{\Gamma\mathbf{v}}\right)^{-1}$
We apply a causal attention mask to $\mathcal{P}^{Z}_{\mathbf{qk}},\mathcal{P}^{Z}_{\mathbf{v}}$, masking out all $j>i$.
@@
$\mathbf{W}[i,j]=\left(1+\sum_{k}\mathcal{P}^{Z}_{\mathbf{qk}}[k,i,j]\cdot\big|\mathcal{R}_{\mathbf{qk}}[k,i,j]\big|^{2}/\nu\right)^{-1}$
@@
$\tilde{\mathcal{A}}[k,i,j]=\mathbf{W}[i,j]\cdot\mathcal{P}_{\mathbf{v}}^{Z}[k,i,j]$
@@ 3.2 Complexity
Standard attention with input length $m$ and embedding dimension $d$ has time complexity $\mathcal{O}(m^{2}d)$ to compute attention scores and apply them across the sequence, plus $\mathcal{O}(md^{2})$ to transform the inputs into query, key, and value vectors and to transform the outputs through a linear layer. The storage of the query, key, and value matrices requires $\mathcal{O}(md)$ memory, and the attention matrix requires $\mathcal{O}(m^{2})$ memory.
The naive tensor implementation of attention above involves element-wise multiplication of $d\times m\times m$ tensors, resulting in time complexity $\mathcal{O}(m^{2}d+md^{2})$, which is the same as standard attention, but slower in practice due to less-optimized element-wise operations. The time complexity to propagate the precision matrices $\mathbf{P}_{ij}^{C}$ through the differential Lyapunov equation in the full (non-diagonalizable case) is $\mathcal{O}(d^{3})$, plus another $\mathcal{O}(d^{3})$ to invert each of these matrices, for each of the $\mathcal{O}(m^{2})$ time differences. Hence, the total time complexity is $\mathcal{O}(m^{2}d^{3})$. In contrast, for systems that admit a closed-form solution, the propagation and matrix inversions become $\mathcal{O}(d)$, so the total time complexity is $\mathcal{O}(m^{2}d)$. If the measurements are sampled at equal time intervals, there are only $\mathcal{O}(m)$ unique time intervals, so the time complexity to compute and invert the precision matrices reduces to $\mathcal{O}(md)$.
@@ 3.3 Exploiting the Toeplitz Structure
$\mathcal{K}_{\mathbf{qk}}^{V}[\tau]=\bm{\lambda}_{C\mathbf{qk}}^{2}\odot\bm{\lambda}_{Q\mathbf{qk}}\odot\frac{\mathbf{1}-e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})\Delta t\tau}}{-2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})}+\bm{\lambda}_{R\mathbf{k}}\odot e^{2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})\Delta t\tau}+\bm{\lambda}_{R\mathbf{q}}$
@@
$\mathcal{K}_{\mathbf{qk}}^{V}[\tau]=\bm{\lambda}_{C\mathbf{qk}}^{2}\odot\bm{\lambda}_{Q\mathbf{qk}}\odot\frac{\mathbf{1}-\mathcal{K}^{\dagger}_{\mathbf{k}}[\tau]\mathcal{K}_{\mathbf{qk}}[\tau]}{-2\mathrm{Re}(\bm{\lambda}_{\mathbf{qk}})}+\bm{\lambda}_{R\mathbf{k}}\odot\big(\mathcal{K}^{\dagger}_{\mathbf{qk}}[\tau]\mathcal{K}_{\mathbf{qk}}[\tau]\big)+\bm{\lambda}_{R\mathbf{q}}$
@@
$\mathcal{P}_{\mathbf{k}}^{Z}[k,i,j]=\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|],$
where $\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|]=\mathcal{K}_{\mathbf{qk}}^{V}[k,|i-j|]^{-1}$, and similarly for $\mathcal{K}_{\mathbf{v}}^{P}[\tau]$ and $\mathcal{P}_{\mathbf{v}}^{Z}$.
@@
3.4 Factorizing the Pulled-Forward Estimates
Because the dynamics are LTI, we can avoid explicitly constructing the pulled-forward estimates $\mathcal{\hat{Z}}_{\mathbf{k}}[k,i,j]$ and $\mathcal{\hat{Z}}_{\mathbf{v}}[k,i,j]$ by defining separate arrays for the $i$ and $j$ components:
@@ 3.5 Expanding the Squared Residuals
$\mathbf{D}[i,j]:=\sum_{k}\mathcal{R}^{*}_{\mathbf{qk}}[k,i,j]\cdot\mathcal{P}^{Z}_{\mathbf{qk}}[k,i,j]\cdot\mathcal{R}_{\mathbf{qk}}[k,i,j]$
@@
$\sum_{k}\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\left(\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]-\mathbf{\hat{Z}}^{*}_{\mathbf{k}}[k,j]\right)\cdot\mathcal{P}^{Z}_{\mathbf{qk}}[k,i,j]\cdot\left(\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]-\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)$
@@
$\mathbf{D}[i,j]=\sum_{k}\mathcal{K}^{P}_{\mathbf{qk}}[k,|i-j|]\cdot\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\cdot\Big[\big|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\big|^{2}+\big|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big|^{2}-2\mathrm{Re}\big(\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big)\Big]$
@@
3.6 Key Simplification: Isotropic Decay and Noise
@@
This corresponds to a system with an isotropic plus skew-symmetric state matrix, $\mathbf{A}=\alpha\mathbf{I}+\bm{\Omega}$, where $\bm{\Omega}\in\mathbb{R}^{d\times d}$ is skew-symmetric ($\bm{\Omega}^{\top}=-\bm{\Omega}$) and diagonalizable over $\mathbb{C}$. That is, $\bm{\Omega}=\mathbf{S}\bm{\Lambda}_{\Omega}\mathbf{S}^{-1}$, where $\mathbf{S}\in\mathbb{C}^{d\times d}$, $\bm{\Lambda}_{\Omega}=\text{diag}(i\omega_{1},\dots i\omega_{d})$, with $\omega_{k}\in\mathbb{R}$.
In general, we can define separate dynamics parameters $\alpha_{\mathbf{qk}},\bm{\omega}_{\mathbf{qk}}$ and $\alpha_{\mathbf{v}},\bm{\omega}_{\mathbf{v}}$ for the queries/keys and the values.
We also assume that the noise is isotropic, i.e. that the noise parameters are scalar multiples of identity: $\bm{\Lambda}_{Q,\mathbf{qk}}=\sigma_{\mathbf{qk}}^{2}\mathbf{I},\,\bm{\Lambda}_{Q,\mathbf{v}}=\sigma_{\mathbf{v}}^{2}\mathbf{I},\,\bm{\Lambda}_{R,\mathbf{k}}=\eta^{2}_{\mathbf{k}}\mathbf{I},\,\bm{\Lambda}_{R,\mathbf{q}}=\eta^{2}_{\mathbf{q}}\mathbf{I},\,\bm{\Lambda}_{R,\mathbf{v}}=\eta^{2}_{\mathbf{v}}\mathbf{I},\,\bm{\Lambda}_{\Gamma,\mathbf{v}}=\gamma^{2}_{\mathbf{v}}\mathbf{I}$.
Alternatively, we can define shared dynamics parameters $\alpha,\bm{\omega}$ and noise parameters $\sigma^{2},\eta^{2},\gamma^{2}$.
Our covariance kernels become:
$\sigma_{V,\mathbf{qk}}^{2}(\tau)=\big|\lambda_{C,\mathbf{qk}}\big|^{2}\sigma_{\mathbf{qk}}^{2}\frac{1-e^{-2\alpha_{\mathbf{qk}}\Delta t\tau}}{-2\alpha_{\mathbf{qk}}}+\eta_{\mathbf{k}}^{2}e^{-2\alpha_{\mathbf{qk}}\Delta t\tau}+\eta_{\mathbf{q}}^{2}$
$\sigma_{V,\mathbf{v}}^{2}(\tau)=\big|\lambda_{C,\mathbf{v}}\big|^{2}\sigma_{\mathbf{v}}^{2}\frac{1-e^{-2\alpha_{\mathbf{v}}\Delta t\tau}}{-2\alpha_{\mathbf{v}}}+\eta_{\mathbf{v}}^{2}e^{-2\alpha_{\mathbf{v}}\Delta t\tau}+\gamma_{\mathbf{v}}^{2}$
@@
$\mathbf{P}_{\mathbf{qk}}^{Z}[i,j]=\frac{1}{\sigma_{V,\mathbf{qk}}^{2}(|i-j|)}.$
We can then pull $\mathbf{P}^{Z}_{\mathbf{qk}}\in\mathbb{R}^{m\times m}$ outside the sum:
$\mathbf{D}[i,j]=\mathbf{P}^{Z}_{\mathbf{qk}}[i,j]\sum_{k}\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\Big[\big|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\big|^{2}+\big|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big|^{2}-2\mathrm{Re}\left(\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)\Big]$
@@
Letting $\mathbf{P}_{\mathbf{v}}^{Z}[i,j]=\frac{1}{\sigma_{V,\mathbf{v}}^{2}(|i-j|)}$ and defining a score matrix $\mathbf{L}=\log(\mathbf{P}_{\mathbf{v}}^{Z})-\log\left(\mathbf{I}+\mathbf{P}^{Z}_{\mathbf{qk}}\odot\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}/\nu\right)$, and a causal mask $\mathbf{M}_{\text{causal}}$, we can then express the row-normalization using softmax:
@@
Note that for $\mathrm{Re(}\bm{\lambda}_{\mathbf{v}})\leq\mathbf{0}$, $\mathbf{\hat{Z}_{v}}$ grows exponentially and $\bm{\Phi}_{\mathbf{v}}$ decays exponentially. However, under the assumption of isotropic decay $\mathrm{Re}(\bm{\lambda}_{\mathbf{v}})=\alpha\mathbf{1}$, we can re-factorize $\mathbf{\bar{Z}_{v}}[k,i]$ such that this problem is avoided.
@@
$\mathbf{Z}_{\mathbf{v}}^{+}=e^{\alpha\Delta t}e^{\bm{\omega}\Delta t}\odot\mathbf{\bar{Z}_{v}}$
And the output is:
$\mathbf{Z}^{+}=\mathbf{W_{p}}\mathbf{Z}_{\mathbf{v}}^{+}$
Intuitively, we rotate the queries, keys, and values into the same frame to perform the attention computation, and then reverse the rotation of the values to return to the original frame.
The assumption that the real part of the eigenvalues is a shared scalar serves two purposes: it allows the cross term to be represented as a matrix multiplication, which reduces the memory requirement of the algorithm from $\mathcal{O}(m^{2}d)$ to $\mathcal{O}(m^{2}+md)$, as in standard attention, while maintaining the $\mathcal{O}(m^{2}d)$ compute time; (b) it improves the exploding/vanishing gradient problem for $\alpha<0$, because rather than computing $e^{\alpha t_{i}}$ and $e^{-\alpha t_{j}}$ separately, we compute $e^{\alpha(t_{i}-t_{j})}$, which is stable, since $t_{i}>t_{j}$ for causal systems. This comes at the price of losing the second-order information provided by the full tensors $\mathcal{P}^{Z}_{\mathbf{qk}}$ and $\mathcal{P}^{Z}_{\mathbf{v}}$. Whereas $\mathcal{P}^{Z}_{\mathbf{qk}},\mathcal{P}^{Z}_{\mathbf{v}}$ represent the confidence of the $k$-th feature’s projection from time $t_{j}$ to time $t_{i}$, $\mathbf{P}^{Z}_{\mathbf{qk}},\mathbf{P}^{Z}_{\mathbf{v}}$ represent measurements of overall confidence between query and key.
Compared to dot-product attention’s single matrix multiplication, computing the residual matrix $\big|\mathbf{R}_{\mathbf{qk}}\big|^{2}$ requires additional vector summations, broadcasting, and element-wise multiplications. However, computing $|\mathbf{Z_{q}|}$ and $|\mathbf{Z_{k}|}$ is only $\mathcal{O}(d_{qk}m)$, so this is a negligible overhead compared to the $\mathcal{O}(d_{qk}m^{2})$ dot-product computation.
3.7 Further Simplification: Normalized Queries & Keys